# Language Models as Semantic Teachers: Post-Training Alignment for Medical Audio Understanding

### v1

@@ IV-C Downstream Tasks and Evaluation Protocol
Downstream Tasks. We evaluate all models on a challenging benchmark of 18 downstream tasks (figure 2B), primarily sourced from the OPERA benchmark [zhang2024towards] and expanded with additional cardiac sound datasets. As detailed in Table II, these tasks cover three distinct clinical areas: respiratory health classification, lung function regression, and cardiac condition classification.
@@
The prediction head is a simple shallow network, either a single linear layer or a one-hidden-layer MLP. Its architecture is selected as a hyperparameter for each task using the same settings as OPERA [zhang2024towards] for consistency. It is trained using the Adam optimizer with an initial learning rate of $10^{-4}$ and an $L_{2}$ penalty. The learning rate is decayed by a factor of $0.97$ after each epoch. We employ an early stopping criterion, halting training if the validation metric fails to improve for five consecutive epochs and retaining the checkpoint with the best validation performance for testing.
@@
Zero‑shot classification. In addition to the linear probe evaluation above, we assess AcuLa in a fully zero‑shot regime (Table VI). For each test clip we (i) extract its frozen embedding, (ii) retrieve the top‑5 clinical reports from the FAISS [douze2024faiss] text index built on the train set embeddings, (iii) embed the retrieved report together with the task’s class names using JINA (text model) [günther2025jinaembeddingsv4universalembeddingsmultimodal], and (iv) assign the class whose text embedding has the highest cosine similarity to the report embedding. No task‑specific weights are learned; the entire pipeline uses our alignment model.
V Results
TABLE V: Comparison of audio encoders after alignment with MedGemma-4B [OPERA, CLAP, AudioMAE] and Qwen 2.5-Omni-7B. T1–T9: respiratory classification [AUROC ($\uparrow$), higher = better]. T10–T16: lung-function estimation [MAE ($\downarrow$), lower = better]. Numbers in brackets give absolute changes vs. the baseline of the same backbone. Improvements are highlighted in green. ++ indicates AcuLa enhanced semantic alignment improves upon pre-trained encoders.
@@ V Results
TABLE VI: AUROC ($\uparrow$) on nine respiratory-classification tasks. Columns use the abbreviations Exh. (exhalation), Cgh. (cough), and COPD sev. (COPD severity). Baselines (left block) are reported in [zhang2024towards]. AcuLa (Zero‑shot): our retrieval‑and‑similarity pipeline that classifies each test clip without seeing any task labels. AcuLa: a task‑specific logistic‑regression probe trained on frozen AcuLa embeddings, following the linear probing protocol described in Section IV-C.
@@ V-A Performance on Downstream Clinical Tasks
We evaluate AcuLa against a suite of strong baselines on 18 downstream tasks. The results, summarized in Tables III and IV, show that AcuLa achieves state-of-the-art performance across respiratory classification, lung function regression, and cardiac classification tasks.
@@
Respiratory Health Condition Classification. In the nine classification tasks (Table III), AcuLa demonstrates superior performance, showing a clear advantage in identifying pathological conditions from audio signals. We see highest improvements in tasks that require nuanced acoustic discrimination. For instance, in COVID-19 detection from cough sounds (T3), AcuLa improves the AUROC to 0.887, a substantial gain over the next-best baseline and far exceeding traditional methods like OpenSMILE (0.537 AUROC). This suggests that the semantic guidance from the LLM helps the model distinguish subtle, clinically significant variations in coughs that purely acoustic models miss. Similarly, major gains in smoker identification (T6) and COPD-related tasks (T5, T8, T9) indicate that AcuLa effectively learns to associate specific acoustic biomarkers with their underlying clinical labels. Even on tasks like gender classification (T4, T7), where acoustic cues are already strong, AcuLa maintains a competitive edge, confirming that the semantic alignment enhances, rather than compromises, the model’s inherent discriminative power.
@@
Lung Function Regression. AcuLa also sets a new standard in all seven lung function estimation tasks, achieving the lowest Mean Absolute Error (MAE) in every case (Table IV). The improvements are particularly strong for tasks involving sustained phonation (T13-T15), where a semantic understanding of vocal effort and respiratory capacity is most beneficial. Here, AcuLa substantially reduces prediction errors for FVC and FEV1, likely because the LLM’s knowledge helps the model interpret how vocal patterns correlate with physiological lung parameters. While the gains on breath-based spirometry tasks (T10-T12) are more modest, they are consistent across all metrics, showing the broad applicability of our approach. The improved accuracy in breathing rate estimation (T16) further underscores AcuLa’s enhanced ability to extract physiologically meaningful information from complex respiratory sounds.
@@
TABLE VII: Average performance across task categories for different LLMs. Respiratory-condition inference uses AUROC $\uparrow$ (higher = better), lung-function estimation uses MAE $\downarrow$ (lower = better), and cardiac-condition inference uses AUROC $\uparrow$ (higher = better). Detailed per-task results are provided in Appendix E, Table XII.
Task Category # Llama-3.2-3B[llama2023] DS-R1-DQ-1.5B[guo2025deepseek] Helium 2B[kyutai_helium1_2025] SmolLM2[allal2024SmolLM2] MedGemma-4B[google_medgemma_2025]
@@
Generality and Model-Agnosticism A core contribution of our work is that AcuLa is a general framework applicable to any pre-trained audio encoder. To validate this, we apply our post-training alignment procedure to a diverse set of encoders, including the OPERA family, CLAP, AudioMAE, and the audio encoder of a general multimodal model, Qwen2.5-Omni[qwen]. The results, presented in Table V, are unequivocal. Regardless of the underlying architecture or original training paradigm, our alignment method consistently and significantly boosts performance. For example, OPERA-based models see large gains in cough-based tasks, while the general-purpose AudioMAE and CLAP models become much more effective at clinical classification after alignment. Even when applied to the audio encoder from Qwen2.5-Omni[qwen], which was not pre-trained on medical data, AcuLa yields competitive results, demonstrating its power to instill domain-specific semantics. This strong and consistent improvement across various backbones confirms that AcuLa is a versatile and model-agnostic framework for enhancing clinical audio understanding.
@@
Zero‑shot respiratory classification. Table VI shows that AcuLa’s zero‑shot pipeline performs competitively across nine respiratory tasks, often rivaling or exceeding the best audio-only baselines (VGGish, AudioMAE, CLAP, OCT, OCE, OGT), despite using no task-specific labels. In several tasks (e.g., Smoker, Covid), the retrieval-based approach demonstrates clear advantages. Further adding a lightweight linear probe to the frozen AcuLa features yields an additional 4–12 pp absolute AUROC gain, leading to new state-of-the-art results on most tasks. See Appendix D for examples of the retrieval outputs.
V-B Ablation Studies and Analysis
Choice of Semantic Teacher (LLM).
First, we investigate how the choice and size of the LLM teacher affect performance. Table VII shows a comparison of five different language models. The domain-specialized MedGemma-4B model consistently provides the best results across all three task categories. This highlights the value of leveraging an LLM with deep, pre-existing medical knowledge, particularly for nuanced classification tasks where it achieves a mean AUROC of 0.786. Interestingly, while smaller, general-purpose models are less effective, their performance on regression tasks suggests that basic linguistic competence is sufficient for capturing some physiologically relevant correlations. However, for high-stakes clinical tasks, a domain-aware teacher is clearly superior than general LLMs.
Domain-Specific Training Table LABEL:tab:resp_only examines the impact of training exclusively on respiratory sounds. While respiratory task performance marginally improves (0.788 vs 0.786 AUROC), cardiac performance degrades (0.601 vs 0.661 AUROC). This trade-off demonstrates that diverse acoustic training enables better cross-domain generalization, even when the primary application domain is well-represented in the training data. Per-task results for this respiratory-only adaptation are reported in Appendix F.
Impact of Training Data. Next, we analyze the influence of the (alignment) training data’s diversity and augmentation, with results summarized in Table LABEL:tab:avg_performance_aug. When training exclusively on respiratory sounds, performance on respiratory tasks sees a negligible change (0.786 vs. 0.788 AUROC), but performance on out-of-domain cardiac tasks drops significantly (0.661 vs. 0.601 AUROC). This demonstrates that training on a large and diverse acoustic corpus is crucial for building a robust, generalizable model, even when the target application is domain-specific. Removing data augmentation reveals its critical role, especially for regression. While classification performance sees a modest decline, the MAE for lung function estimation increases dramatically from 0.821 to 0.973. This suggests that augmentation forces the model to learn features that are invariant to superficial recording variations (like volume), which is essential for predicting precise continuous values. A full per-task comparison is provided in Appendix G.
Alignment Strategy. We compare two alignment strategies: aligning only the final transformer layer (Last-L) versus multiple intermediate layers (Multi-B). As shown in Table LABEL:tab:align_avg, the simpler single-layer approach is generally superior or comparable. The multi-layer approach offers no consistent benefit, indicating that the final layer of the LLMs already contains a sufficiently rich and compressed representation for semantic alignment. A task-wise breakdown is given in Appendix H.
@@ V-B Ablation Studies and Analysis
TABLE VIII: Summary of ablation studies. Each table reports the average performance over the three task categories. Respiratory and cardiac tasks use AUROC $\uparrow$ (higher = better); lung-function tasks use MAE $\downarrow$ (lower = better).
@@
TABLE IX: Ablation on masked–reconstruction loss (Mask-Rec) and alignment loss (Align) with a pre-trained model. The fourth row swaps our alignment loss (i.e., CKA-based) for an $\ell^{2}$ (MSE) loss (marked $\dagger$). The first row (no check-mark in “Pre-trained”) represents randomly initialized audio encoder but applies both losses.
Mask-Rec Align Pre-trained Respiratory health $\uparrow$ Lung function $\downarrow$ Cardiac condition $\uparrow$
@@
Loss Component Analysis. We ablate the contributions of our dual-objective function components. As shown in Table IX, both the masked reconstruction loss (Mask-Rec) and alignment loss (Align) contribute to final performance, with the alignment loss proving more critical.
Removing masked reconstruction while keeping alignment results in modest degradation: respiratory condition inference drops from 0.786 to 0.768 AUROC, lung function estimation worsens from 0.821 to 0.865 MAE, and cardiac inference decreases from 0.661 to 0.645 AUROC. In contrast, removing alignment loss while keeping masked reconstruction causes larger performance drops across all categories: respiratory (0.786 to 0.715), lung function (0.821 to 0.904 MAE), and cardiac (0.661 to 0.591).
Furthermore, we examine alternative alignment objectives. Using $\ell^{2}$ (MSE) loss instead of CKA for alignment (marked †) results in slightly worse performance across all metrics, validating our choice of CKA. Finally, training from random weights instead of a pre-trained model catastrophically degrades performance (last row).
These results demonstrate that the CKA-based alignment loss is the primary driver for instilling semantic understanding, while masked reconstruction acts as a valuable regularizer that prevents the audio encoder from forgetting its acoustic modeling capabilities during alignment.
@@ VI Conclusion
In this work, we demonstrate that pre-trained large language models can serve as effective “semantic teachers” to instill deep clinical understanding into specialized audio encoders. We introduced AcuLa, a general, lightweight alignment framework that successfully grounds high-level medical concepts from text into the fine-grained, temporal patterns of cardio-respiratory sounds. Our comprehensive experiments show that this fusion of semantic knowledge and acoustic modeling creates representations that are not only superior across a diverse range of 18 classification and regression tasks but are also more robust and clinically relevant. Our work establishes a novel direction for cross-modal learning, inverting the traditional knowledge flow to enhance perceptual models with abstract semantics. While our data generation strategy offers a scalable solution to leverage metadata for clinical text scarcity problem, the true promise lies in deploying this paradigm in data-rich clinical environments. Future work could extend this teacher-student paradigm to other physiological time-series like EEG and ECG, or develop self-correction cycles where model-disagreements flag cases for human-in-the-loop review, moving towards AI systems that truly reason about clinical data.
@@ Acknowledgments
This work was supported by the NWO AiNed Fellowship Grant of A.S. We also acknowledge the use of the Dutch National Supercomputer Snellius for essential computational tasks.

### v2

@@ IV-C Downstream Tasks and Evaluation Protocol
Downstream Tasks. We evaluate all models on a challenging benchmark of 18 downstream tasks (figure 1B), primarily sourced from the OPERA benchmark [32] and expanded with additional cardiac sound datasets. As detailed in Table III, these tasks cover three distinct clinical areas: respiratory health classification, lung function regression, and cardiac condition classification.
@@
The prediction head is a simple shallow network, either a single linear layer or a one-hidden-layer MLP. Its architecture is selected as a hyperparameter for each task using the same settings as OPERA [32] for consistency. It is trained using the Adam optimizer with an initial learning rate of $10^{-4}$ and an $L_{2}$ penalty. The learning rate is decayed by a factor of $0.97$ after each epoch. We employ an early stopping criterion, halting training if the validation metric fails to improve for five consecutive epochs and retaining the checkpoint with the best validation performance for testing.
@@
Zero‑shot classification. In addition to the linear probe evaluation above, we assess AcuLa in a fully zero‑shot regime (Table VII). For each test clip we (i) extract its frozen embedding, (ii) retrieve the top‑5 clinical reports from the FAISS [4] text index built on the train set embeddings, (iii) embed the retrieved report together with the task’s class names using JINA (text model) [günther2025jinaembeddingsv4universalembeddingsmultimodal], and (iv) assign the class whose text embedding has the highest cosine similarity to the report embedding. No task‑specific weights are learned; the entire pipeline uses our alignment model.
TABLE VI: Comparison of audio encoders after alignment with MedGemma-4B [OPERA, CLAP, AudioMAE] and Qwen 2.5-Omni-7B. T1–T9: respiratory classification [AUROC ($\uparrow$), higher = better]. T10–T16: lung-function estimation [MAE ($\downarrow$), lower = better]. Numbers in brackets give absolute changes vs. the baseline of the same backbone. Improvements are highlighted in green. ++ indicates AcuLa enhanced semantic alignment improves upon pre-trained encoders.
@@
V Results
TABLE VII: AUROC ($\uparrow$) on nine respiratory-classification tasks. Columns use the abbreviations Exh. (exhalation), Cgh. (cough), and COPD sev. (COPD severity). Baselines (left block) are reported in [32]. AcuLa (Zero‑shot): our retrieval‑and‑similarity pipeline that classifies each test clip without seeing any task labels. AcuLa: a task‑specific logistic‑regression probe trained on frozen AcuLa embeddings, following the linear probing protocol described in Section IV-C.
@@ V-A Performance on Downstream Clinical Tasks
We evaluate AcuLa against a suite of strong baselines on 18 downstream tasks. The results, summarized in Tables IV and V, show that AcuLa achieves state-of-the-art performance across respiratory classification, lung function regression, and cardiac classification tasks.
@@
Respiratory Health Condition Classification. In the nine classification tasks (Table IV), AcuLa demonstrates superior performance, showing a clear advantage in identifying pathological conditions from audio signals. We see highest improvements in tasks that require nuanced acoustic discrimination. For instance, in COVID-19 detection from cough sounds (T3), AcuLa improves the AUROC to 0.887, a substantial gain over the next-best baseline and far exceeding traditional methods like OpenSMILE (0.537 AUROC). This suggests that the semantic guidance from the LLM helps the model distinguish subtle, clinically significant variations in coughs that purely acoustic models miss. Similarly, major gains in smoker identification (T6) and COPD-related tasks (T5, T8, T9) indicate that AcuLa effectively learns to associate specific acoustic biomarkers with their underlying clinical labels. Even on tasks like gender classification (T4, T7), where acoustic cues are already strong, AcuLa maintains a competitive edge, confirming that the semantic alignment enhances, rather than compromises, the model’s inherent discriminative power.
@@
Lung Function Regression. AcuLa also sets a new standard in all seven lung function estimation tasks, achieving the lowest Mean Absolute Error (MAE) in every case (Table V). The improvements are particularly strong for tasks involving sustained phonation (T13-T15), where a semantic understanding of vocal effort and respiratory capacity is most beneficial. Here, AcuLa substantially reduces prediction errors for FVC and FEV1, likely because the LLM’s knowledge helps the model interpret how vocal patterns correlate with physiological lung parameters. While the gains on breath-based spirometry tasks (T10-T12) are more modest, they are consistent across all metrics, showing the broad applicability of our approach. The improved accuracy in breathing rate estimation (T16) further underscores AcuLa’s enhanced ability to extract physiologically meaningful information from complex respiratory sounds.
Fair in-domain baseline adaptation. To separate gains from semantic alignment vs. in-domain audio exposure, we fine-tune AudioMAE and CLAP on the same alignment training audio (train splits only) without reports, and evaluate all models using the identical frozen linear-probe protocol. Audio-only adaptation improves both baselines, yet AcuLa remains consistently stronger overall, especially on respiratory and cardiac inference, which indicates benefits beyond domain adaptation alone. Table IX summarizes category-level results; per-task results are in Appendix F (Tables XVIIXIX).
@@
TABLE VIII: Average performance across task categories for different LLMs. Respiratory-condition inference uses AUROC $\uparrow$ (higher = better), lung-function estimation uses MAE $\downarrow$ (lower = better), and cardiac-condition inference uses AUROC $\uparrow$ (higher = better). Detailed per-task results are provided in Appendix D, Table XIV.
Task Category # Llama-3.2-3B[29] DS-R1-DQ-1.5B[11] Helium 2B[18] SmolLM2[1] MedGemma-4B[10]
@@
TABLE IX: Fair in-domain baseline adaptation (train splits only, no reports), evaluated with the same frozen linear-probe protocol. Resp.=T1–T9 AUROC $\uparrow$, LungFn.=T10–T16 MAE $\downarrow$, Cardiac=T17–T18 AUROC $\uparrow$.
Group AudioMAE-ft CLAP-ft AcuLa
Resp. 0.686 0.723 0.786
LungFn. 0.864 0.871 0.821
Cardiac 0.629 0.637 0.661
@@
Generality and Model-Agnosticism A core contribution of our work is that AcuLa is a general framework applicable to any pre-trained audio encoder. To validate this, we apply our post-training alignment procedure to a diverse set of encoders, including the OPERA family, CLAP, AudioMAE, and the audio encoder of a general multimodal model, Qwen2.5-Omni[6]. The results, presented in Table VI, are unequivocal. Regardless of the underlying architecture or original training paradigm, our alignment method consistently and significantly boosts performance. For example, OPERA-based models see large gains in cough-based tasks, while the general-purpose AudioMAE and CLAP models become much more effective at clinical classification after alignment. Even when applied to the audio encoder from Qwen2.5-Omni[6], which was not pre-trained on medical data, AcuLa yields competitive results, demonstrating its power to instill domain-specific semantics. This strong and consistent improvement across various backbones confirms that AcuLa is a versatile and model-agnostic framework for enhancing clinical audio understanding.
@@
Zero‑shot respiratory classification. Table VII shows that AcuLa’s zero‑shot pipeline performs competitively across nine respiratory tasks, often rivaling or exceeding the best audio-only baselines (VGGish, AudioMAE, CLAP, OCT, OCE, OGT), despite using no task-specific labels. In several tasks (e.g., Smoker, Covid), the retrieval-based approach demonstrates clear advantages. Further adding a lightweight linear probe to the frozen AcuLa features yields an additional 4–12 pp absolute AUROC gain, leading to new state-of-the-art results on most tasks. See Appendix G for examples of the retrieval outputs. We further quantify cross-modal semantic alignment with bidirectional audio–text retrieval on held-out pairs (Appendix H).
@@
TABLE X: Summary of ablation studies. Each table reports the average performance over the three task categories. Respiratory and cardiac tasks use AUROC $\uparrow$ (higher = better); lung-function tasks use MAE $\downarrow$ (lower = better).
@@
TABLE XI: Ablation of masked reconstruction (Mask-Rec) and alignment (Align) losses with a pre-trained model. $\dagger$ replaces our CKA-based alignment with $\ell^{2}$ (MSE). The first row uses a randomly initialized audio encoder with both losses.
Mask-Rec Align Pre-trained Resp. health $\uparrow$ Lung func. $\downarrow$ Cardiac cond. $\uparrow$
@@
V-B Ablation Studies and Analysis
Choice of Semantic Teacher (LLM). First, we examine how LLM teacher choice and size affect performance. As shown in Table VIII, the domain-specialized MedGemma-4B performs best across all three task categories. This underscores the benefit of strong medical prior knowledge, especially for nuanced classification, where it achieves a mean AUROC of 0.786. Smaller general-purpose models perform worse overall, though their regression results suggest that basic linguistic competence can still capture some physiologically relevant correlations. For high-stakes clinical tasks, however, a domain-aware teacher is clearly more effective.
Domain-Specific Training Table LABEL:tab:resp_only examines the impact of training exclusively on respiratory sounds. While respiratory task performance marginally improves (0.788 vs 0.786 AUROC), cardiac performance degrades (0.601 vs 0.661 AUROC). This trade-off demonstrates that diverse acoustic training enables better cross-domain generalization, even when the primary application domain is well-represented in the training data. Per-task results for this respiratory-only adaptation are reported in Appendix I.
Impact of Training Data. We analyze how (alignment) training data diversity and augmentation affect performance (Table LABEL:tab:avg_performance_aug). Training only on respiratory sounds leaves respiratory tasks nearly unchanged (0.786 vs. 0.788 AUROC) but substantially degrades out-of-domain cardiac performance (0.661 vs. 0.601 AUROC), highlighting the importance of large, diverse acoustic corpora for generalization. Removing augmentation has a modest impact on classification but severely hurts regression: lung-function MAE increases from 0.821 to 0.973, suggesting augmentation promotes invariance to superficial recording variations (e.g., volume). Per-task results are in Appendix J.
Alignment Strategy. We compare two alignment strategies: aligning only the final transformer layer (Last-L) versus multiple intermediate layers (Multi-B). As shown in Table LABEL:tab:align_avg, the simpler single-layer approach is generally superior or comparable. The multi-layer approach offers no consistent benefit, indicating that the final layer of the LLMs already contains a sufficiently rich and compressed representation for semantic alignment. A task-wise breakdown is given in Appendix K.
Loss Component Analysis. We ablate the two components of our dual-objective loss. As shown in Table XI, both masked reconstruction (Mask-Rec) and alignment (Align) improve performance, with alignment being more important. Removing Mask-Rec causes modest drops: respiratory AUROC falls from 0.786 to 0.768, lung function MAE rises from 0.821 to 0.865, and cardiac AUROC declines from 0.661 to 0.645. Removing Align leads to much larger degradations across all tasks: respiratory drops to 0.715, lung MAE worsens to 0.904, and cardiac falls to 0.591. We also test alternative alignment objectives. Replacing CKA with $\ell^{2}$ (MSE) loss gives slightly worse results across all metrics, supporting our choice of CKA. Training from random initialization instead of a pre-trained model severely hurts performance. Overall, these results show that CKA-based alignment is the main source of semantic learning, while masked reconstruction serves as a useful regularizer that preserves acoustic modeling ability during alignment.
Label-masked report ablation. To assess whether AcuLa’s gains are driven by explicit diagnostic keywords, we performed a label-masked ablation in which label-revealing fields/words were removed during report generation and the model was re-trained using audio paired with the masked reports. As shown in Table XVI and detailed in Appendix E, this causes only minor performance changes, with the largest effect on overlap tasks and negligible change elsewhere. These results suggest that AcuLa’s gains are not primarily driven by trivial label-text leakage.
@@ VI Conclusion
In this work, we demonstrate that pre-trained large language models can serve as effective “semantic teachers” to inject clinically meaningful semantic structure into specialized audio encoders. We introduced AcuLa, a general, lightweight alignment framework that successfully grounds high-level medical concepts from text into the fine-grained, temporal patterns of cardio-respiratory sounds. Our comprehensive experiments show that this fusion of semantic knowledge and acoustic modeling creates representations that are not only superior across a diverse range of 18 classification and regression tasks but are also more robust and clinically relevant. Our work establishes a novel direction for cross-modal learning, inverting the traditional knowledge flow to enhance perceptual models with abstract semantics. While our data generation strategy offers a scalable solution to leverage metadata for clinical text scarcity problem, the true promise lies in deploying this paradigm in data-rich clinical environments. Future work could extend this teacher-student paradigm to other physiological time-series like EEG and ECG, or develop self-correction cycles where model-disagreements flag cases for human-in-the-loop review, moving towards AI systems that truly reason about clinical data.
@@ Acknowledgments
This work was supported by the NWO AiNed Fellowship Grant awarded of A.S., and in part by Google.org and the Google Cloud Research Credits program through the Gemini Academic Program. We also acknowledge the use of the Dutch National Supercomputer Snellius for essential computational tasks. We thank Martijn den Dekker of Erasmus University Medical Center for the review and feedback.